# Halo Mass Functions at High Redshift

### v1

@@ Abstract
Recent JWST observations of very early galaxies, at $\rm{z\gtrsim 10}$, has led to claims that tension exists between the sizes and luminosities of high redshift galaxies and what is predicted by standard $\Lambda$CMD models. Here we use the adaptive mesh refinement code Enzo and the N-body smoothed particle hydrodyanmics code SWIFT to compare (semi-)analytic halo mass functions against the results of direct N-body models at high redshift. In particular, our goal is to investigate the variance between standard halo mass functions derived from (semi-)analytic formulations and N-body calculations and to determine what role any discrepancy may play in driving tensions between observations and theory. We find that the difference between direct N-body calculations and halo mass function fits is less than a factor of two within the mass range of galaxies currently being observed by JWST and is therefore not a dominant source of error when comparing theory and observation at high redshift.
@@ 1 Introduction
Since its launch in December 2021 and its subsequent data releases in the months since, JWST has both reshaped and challenged our understanding of galaxy formation in the very early Universe ($z\gtrsim 10$). In particular, JWST has discovered a number of very massive and very luminous galaxies at redshifts in excess of $z=10$ which after initial photometric detection have now been spectroscopically confirmed (Harikane et al., 2024; Arrabal Haro et al., 2023; Castellano et al., 2024; Hainline et al., 2024; Gentile et al., 2024). These galaxies, both in terms of their intrinsic luminosity and their potential host halo masses, provide a significant challenge to our understanding of structure formation in the early Universe. Boylan-Kolchin (2023) using an analytic model applied to the inferred JWST stellar masses of a number of high redshift source from Labbé et al. (2023) found that the stellar masses implied by the sources required star formation efficiencies, $\epsilon$, significantly in excess of those from the present day universe and perhaps as implausibly high as $\epsilon=1$. The model employed by Boylan-Kolchin (2023) hinges on a number of simple yet strong assumptions. These assumptions use the inferred stellar host mass to derive a host halo mass based on the ratios between the cosmic baryon density and the cosmic matter density and the star formation efficiency. These simple arguments culminate in a calculation of the probability of finding such (luminous) galaxies at early times in a $\Lambda$CDM universe.
@@
Boylan-Kolchin (2023) conclude that the most massive JWST galaxies detected are both at the very limit of galaxy formation theory and that their number densities are difficult to equate with the JWST field of view. They conclude that these issues indicate that there are several unresolved issues in our theories. However, underneath the model employed by Boylan-Kolchin (2023) are a number of astrophysical assumptions which are used when converting the broadband spectral energy distribution to a host stellar mass. This conversion is open to large uncertainties - particularly since our knowledge of how to do this conversion comes from the local Universe. There is no guarantee that these conversion relations can be directly mapped to the high-$z$ Universe and in fact most analysis shows that it is almost certainly not the case (e.g. Kannan et al., 2023; Lu et al., 2024). For example recent analysis by Steinhardt et al. (2023) show that modifying the host population IMF of the inferred stellar population, results in a decrease in the stellar mass by factors of between 10 and 50. Such decreases in the host stellar mass have a direct knock-on effect to the inferred host halo mass and can significantly decrease the tension with $\Lambda$CDM models.
@@
Adding to this important point is that Boylan-Kolchin (2023) (as well as many other studies) utilise the well tested and parameterised halo mass function (HMF) derived by Sheth & Tormen (1999) in order to compute their galaxy number densities and cumulative comoving number density of galaxies. An important consideration therefore is to test, using explicit N-body, calculations how accurate the underlying HMF is at high-$z$ (i.e. at $z\gtrsim 10$), compared to direct N-body calculations, and what error may be associated with this model. This is the goal of this study.
@@
While preparing this work a study by Yung et al. (2024) preformed a similar investigation. Using a high fidelity suite of N-body simulations across a broad range of boxsizes and redshifts Yung et al. (2024) were able to show that HMFs derived from fitting functions and analytical approximations match extremely well to their N-body simulations. Yung et al. (2024) found that even up to $z=15$ the match between their N-body simulations and the fitting functions is no more than approximately a factor of two across a range of fitting functions. As discussed above, providing robust quantification of the match between fitting functions and direct N-body calculations at high-$z$ is extremely timely given the recent JWST results.
@@
Here we perform a similar analysis to Yung et al. (2024) with the difference being that we compare HMF fitting functions against two fundamentally different numerical codes - Enzo (Bryan et al., 2014; Brummel-Smith et al., 2019) and SWIFT (Schaller et al., 2018, 2024). While Yung et al. (2024) used the publicly available Gadget-2 code Springel (2005) which is a TreePM code, with a similar gravitational solver to SWIFT we also use the Particle-Mesh based Enzo code which gives an additional layer of comparison. Our analysis confirms the results of Yung et al. (2024) and we also observe a factor of approximately two difference between the HMFs generated by the N-body codes and the fitting functions.
The structure of the paper is as follows: In §2 we outline the methodology including the simulations and fitting functions employed. In §3 we deliver the results of our analysis and in §4 we summarize and dicuss our results in light of recent JWST observations.
@@ 2.1 Numerical simulations
Table 1: Mass resolutions of each simulation box. All simulations but the $L=7.5$ cMpc/h, $N^{1/3}=1024$ box are run using both Enzo and SWIFT ; this box is run using SWIFT only.
@@
Table 2: Highest space resolutions of each simulation box. All simulations but the $L=7.5$ cMpc/h, $N^{1/3}=1024$ box are run using both Enzo and SWIFT ; this box is run using SWIFT only.
Our simulations begin at $z=127.0$ with initial conditions set using MUSIC (Hahn & Abel (2011)) and end at $z=10.0$. We use a $\Lambda$CDM cosmology with $\text{h}=0.6774$, $\Omega_{\text{m}}=0.2592$, $\sigma_{8}=0.8159$, $n_{\text{eff}}=0.9667$ and the Eisenstein & Hu (1998) transfer function for no baryon acoustic oscillations. The two codes we use, Enzo and SWIFT, are both well tested and have been used extensively within the community. Additionally, both codes use somewhat different strategies for solving the Poisson equation which allows for additional comparison between the various semi-analytic fits against numerical solutions. We now describe both codes but refer the interested reader to the code method papers for more details.
Enzo is a grid-based N-body code with the capability for adapative mesh refinement (AMR), widely used in cosmological hydrodynamics simulations. The AMR allows for improved resolution in areas of interest (e.g. collapsing structures) without greatly increasing computational cost and without needing prior knowledge of the volume to pre-select areas for increased refinement. The gravity solver works by implementing a Fast Fourier Transform (FFT) technique to solve Poisson’s equation at the root grid of each timestep of the simulation. The boundary conditions on the subgrids are then interpolated from the parent grid and the Poisson equation is solved at each time step one subgrid at a time. For a unigrid Enzo simuation (i.e. no refinement), the minimum inter-particle separation in which gravity acts is twice the length of two cells, given as:
$\Delta x_{\text{Unigrid}}=2\frac{L}{N^{1/3}}.$ (1)
When running Enzo simulations, we used a maximum refinement level of 8 meaning that the resolution was increased by a factor 28 in regions of high particle density. This means the highest resolution is given as:
$\Delta x_{\text{Enzo}}=\frac{1}{2^{8}}\Delta x_{\text{Unigrid}}=\frac{1}{2^{7}}\frac{L}{N^{1/3}}.$ (2)
For the dark matter only simulations carried out here refinement is triggered once the particle over-density reaches a factor of 4 greater than the mean density (i.e. MinimumOverDensityForRefinement = 4). In addition to this we set the MinimumMassForRefinementLevelExponent = -0.1 which makes the refinement scheme super-Lagrangian and allows for the higher levels of refinement to be more easily triggered. The combination of these parameter choices mean that our simulation setup is somewhat conservative employing an aggressive refinement strategy (see e.g. O’Shea et al., 2005).
@@
$\Delta x_{\text{SWIFT}}=\frac{1}{5^{2}}\frac{L}{N^{1/3}}.$ (3)
@@ 2.2 Numerical Halo Finders
We analyse the simulation snapshots from our full suite of outputs at $z=20.0$, $z=15.0$ and $z=10.0$ using a friends-of-friends (FOF) (Efstathiou et al. (1985)) and HOP halo (Eisenstein & Hut (1998)) finder (results from the HOP finder can be found in the appendix of the paper as the results are very similar between FOF and HOP and the goal of this study is not to compare halo finders). The FOF halo finder accounts for the distances between dark matter particles within a single snapshot. We use a linking length $\Delta x(b)=b\times L/N^{1/3}$, where $L/N^{1/3}$ is the mean inter-particle separation. This refers to the maximum permitted separation between two particles. A group i.e. halo is found from a set of inter-linked particles.
@@
The HOP halo finder accounts for the distances between particles as well as the computed density of each particle. Rather than creating a continuous density field in the simulation volume, the density of each particle at its given position is estimated using a spherically symmetric cubic spline kernel (Monaghan & Lattanzio (1985)) on its $N_{\text{dens}}=64$ nearest neighbours. The density is normalised to the average density of particles within the simulation box e.g. a particle with a computed density of $\delta=80$ refers to a position 80 times denser than the average i.e. an overdensity (Skory et al. (2010)). A link is made by hopping from a given particle to the densest of its $N_{\text{hop}}=64$ neighbours. This process continues, forming a chain of increasing particle density until we reach a particle that is its own densest neighbour. All chains sharing the same densest particle are part of the same group. If the maximum particle density within a group is above a set threshold overdensity $\delta_{\text{peak}}$, that group is defined as a halo. We use a threshold overdensity of $\delta_{\text{peak}}=100$. This value was chosen to calibrate the HOP finder with the FOF finder so both halo finders found a similar halo count and mass range for the Enzo , $L=0.5$ cMpc/h, $N_{\text{DM}}=512^{3}$, $z=10$ snapshot.
The halo masses found from these halo catalogues are the FOF and HOP masses i.e. the mass of the halo is defined as the sum of the masses of the particles that make up the halos. After we create halo catalogues based on these halo finders, we only include halos made up of at least $N_{\text{DM, min}}=100$ particles to reduce numerical over-counting (i.e. haloes must be well resolved (by at least 100 particles) before we identify them as haloes).
@@ 2.3 Numerical Halo Number Densities
A halo catalogue is created from each snapshot with a specific halo finder using the yt [^1] We use a version of yt called yt-swift developed by Rennehan (2022) as the most recently available version of yt cannot load SWIFT output data. and yt-astro-analysis packages (Turk et al. (2011), Smith et al. (2022)). An array of halo mass values is generated from each halo catalogue, filtered to exclude halos consisting of $<\,N_{\text{DM, min}}$ particles. For each array, the halo mass values are binned on a logarithmic scale into 24 histogram bins with $M_{\text{min}}=10^{3.75}\,\text{M}_{\odot}/\text{h}$ and $M_{\text{max}}=10^{11.75}\,\text{M}_{\odot}/\text{h}$. Halo catalogues from a given simulation suite, halo finder and redshift, $z$, are combined, giving rise to 12 datasets (see Table 3 for more detail). For each Enzo dataset, we combine halo catalogues based on 12 simulation boxes. For each SWIFT dataset, we combine halo catalogues based on 13 simulation boxes (see Tables 1 and 2 for more detail). In our algorithm below, we index datasets with $i\in\{1,\cdots,12\}$, mass bins with $j\in\{1,\cdots,24\}$, Enzo simulation boxes with $k\in\{1,\cdots,12\}$ and SWIFT simulation boxes with $k\in\{1,\cdots,13\}$.
@@
Table 3: The 12 datasets derived by combining halo catalogues for each simulation, halo finder and redshift combination.
@@
We compare our numerical halo number densities with those derived from both analytical forms and popular fits. The number density ($\text{h}^{3}\,\text{cMpc}^{-3}$) of dark matter halos at a given redshift, $z$, in a mass bin centred at $\log_{10}(M_{\text{mid}}/\text{h}^{-1}\,\text{M}_{\odot})$ with a width of $\Delta\log_{10}M$ is defined as:
@@
The halo mass function ($\text{h}^{4}\,\text{cMpc}^{-3}\,\text{M}_{\odot}^{-1}$) is the differential halo number density per unit mass and is defined as:
@@
Here $\rho_{0}$ ($\text{h}^{2}\,\text{M}_{\odot}\,\text{cMpc}^{-3}$) and $\sigma(z,M)$ refer to the mean density of the universe and mass variance respectively. The mass variance is given as:
@@
Figure 1: Comparing the Enzo and SWIFT halo number densities with five halo number densities derived from fits (see Table 4) at $z=20.0$ (upper panel), $z=15.0$ (centre panel) and $z=10.0$ (lower panel) using the FOF halo finder. The black rectangles indicate the regions for more detailed analysis as seen in Figure 2. The yellow shaded regions represent the approximate range of halo masses detected by JWST at $z>10$. Over the halo mass range selected all of the models (both numerical and analytic) agreement at $z=10.0$ is excellent with some deviation at higher redshift.
Figure 2: A zoom-in onto the black rectangles identified in Figure 1. These mass ranges represent the most massive haloes accessible via numerical simulation at these redshifts. At this higher level of detail we see some discrepancy between the numerical halo mass functions of Enzo and SWIFT and their (semi-)analytical counterparts, particularly at $z=15.0$. The residuals for each of the lines is shown in Figure 3 and Figure 4

### v2

@@ Abstract
Recent JWST observations of very early galaxies, at $\rm{z\gtrsim 10}$, have led to claims that tension exists between the sizes and luminosities of high-redshift galaxies and what is predicted by standard $\Lambda$CDM models. Here we use the adaptive mesh refinement code Enzo and the N-body smoothed particle hydrodynamics code SWIFT to compare (semi-)analytic halo mass functions against the results of direct N-body models at high redshift. In particular, our goal is to investigate the variance between standard halo mass functions derived from (semi-)analytic formulations and N-body calculations and to determine what role any discrepancy may play in driving tensions between observations and theory. We find that the difference between direct N-body calculations and (semi-) analytic halo mass function fits is less than a factor of 2 (at $\rm{z\sim 10}$) within the mass range of galaxies currently being observed by JWST, and is therefore not a dominant source of error when comparing theory and observation at high redshift.
@@ 1 Introduction
Since its launch in December 2021 and its subsequent data releases in the months since, JWST has both reshaped and challenged our understanding of galaxy formation in the very early Universe ($z\gtrsim 10$). In particular, JWST has discovered a number of very massive and very luminous galaxies at redshifts in excess of $z=10$ which, after initial photometric detection, have now been spectroscopically confirmed (Harikane et al., 2024; Arrabal Haro et al., 2023a; Castellano et al., 2024; Hainline et al., 2024; Gentile et al., 2024). These galaxies, both in terms of their intrinsic luminosity and their potential host halo masses, provide a significant challenge to our understanding of structure formation in the early Universe. Boylan-Kolchin (2023), using an analytic model applied to the inferred JWST stellar masses of a number of high-redshift sources from Labbé et al. (2023), found that the stellar masses implied by the sources required star formation efficiencies, $\epsilon$, significantly in excess of those from the present-day universe and perhaps as implausibly high as $\epsilon=1$. The model employed by Boylan-Kolchin (2023) hinges on a number of simple yet strong assumptions. These assumptions use the inferred stellar host mass to derive a host halo mass based on the ratios between the cosmic baryon density and the cosmic matter density and the star formation efficiency. These simple arguments culminate in a calculation of the probability of finding such (luminous) galaxies at early times in a $\Lambda$CDM universe.
@@
Boylan-Kolchin (2023) conclude that the most massive JWST galaxies detected are both at the very limit of galaxy formation theory and that their number densities are difficult to equate with the JWST field of view. They conclude that these issues indicate that there are several unresolved issues in our theories. However, underneath the model employed by Boylan-Kolchin (2023) are a number of astrophysical assumptions which are used when converting the broadband spectral energy distribution to a host stellar mass. This conversion is open to large uncertainties - particularly since our knowledge of how to do this conversion comes from the local Universe. There is no guarantee that these conversion relations can be directly mapped to the high-$z$ Universe and in fact most analysis shows that it is almost certainly not the case (e.g. Kannan et al., 2023; Lu et al., 2024). For example, recent analysis by Steinhardt et al. (2023) show that modifying the host population IMF of the inferred stellar population results in a decrease in the stellar mass by factors of between 10 and 50. Such decreases in the host stellar mass have a direct knock-on effect to the inferred host halo mass and can significantly decrease the tension with $\Lambda$CDM models.
@@
Adding to this important point is that Boylan-Kolchin (2023) (as well as many other studies) utilise the well-tested and parameterised halo mass function (HMF) derived by Sheth & Tormen (1999) in order to compute their galaxy number densities and cumulative comoving number density of galaxies. An important consideration therefore is to test, using explicit N-body calculations, how accurate the underlying HMF is at high-$z$ (i.e. at $z\gtrsim 10$), compared to direct N-body calculations, and what error may be associated with this model. This is the goal of this study.
@@
While preparing this work a study by Yung et al. (2024) performed a similar investigation. Using a high-fidelity suite of N-body simulations across a broad range of box sizes and redshifts, Yung et al. (2024) were able to show that HMFs derived from fitting functions and analytical approximations match extremely well to their N-body simulations. Yung et al. (2024) found that even up to $z=15$ the match between their N-body simulations and the fitting functions is no more than approximately a factor of two across a range of fitting functions. As discussed above, providing robust quantification of the match between fitting functions and direct N-body calculations at high-$z$ is extremely timely given the recent JWST results.
@@
Here we perform a similar analysis to Yung et al. (2024), with the difference being that we compare HMF fitting functions against two fundamentally different numerical codes - Enzo (Bryan et al., 2014; Brummel-Smith et al., 2019) and SWIFT (Schaller et al., 2018, 2024). While Yung et al. (2024) used the publicly available Gadget-2 code (Springel (2005)) which is a TreePM code, with a similar gravitational solver to SWIFT , we also use the Particle-Mesh based Enzo code which gives an additional layer of comparison. Our analysis confirms the results of Yung et al. (2024) and we also observe a factor of approximately two difference between the HMFs generated by the N-body codes and the fitting functions.
The structure of the paper is as follows: In §2 we outline the methodology including the simulations and fitting functions employed. In §3 we deliver the results of our analysis and in §4 we summarize and discuss our results in light of recent JWST observations.
@@ 2.1 Numerical simulations
Table 1: Mass resolutions of each simulation box. All simulations but the $L=7.5$ cMpc/h, $N_{\text{DM}}^{1/3}=1024$ box are run using both Enzo and SWIFT ; this box is run using SWIFT only.
@@
Table 2: Highest space resolutions of each simulation box. All simulations but the $L=7.5$ cMpc/h, $N_{\text{DM}}^{1/3}=1024$ box are run using both Enzo and SWIFT ; this box is run using SWIFT only.
Our simulations begin at $z=127.0$ with initial conditions set using MUSIC (Hahn & Abel (2011)) and end at $z=10.0$. We use a $\Lambda$CDM cosmology with $\text{h}=0.6774$, $\Omega_{\text{m},0}=0.2592$, $\sigma_{8}=0.8159$, $n_{\text{eff}}=0.9667$ and the Eisenstein & Hu (1998) transfer function for no baryon acoustic oscillations. The two codes we use, Enzo and SWIFT , are both well-tested and have been used extensively within the community. Additionally, both codes use somewhat different strategies for solving the Poisson equation which allows for additional comparison between the various semi-analytic fits against numerical solutions. We now describe both codes but refer the interested reader to the code method papers for more details.
Enzo is a grid-based N-body code with the capability for adapative mesh refinement (AMR), widely used in cosmological hydrodynamics simulations. The AMR allows for improved resolution in areas of interest (e.g. collapsing structures) without greatly increasing computational cost and without needing prior knowledge of the volume to pre-select areas for increased refinement. The gravity solver works by implementing a Fast Fourier Transform (FFT) technique to solve Poisson’s equation at the root grid of each timestep of the simulation. The boundary conditions on the subgrids are then interpolated from the parent grid and the Poisson equation is solved at each time step, one subgrid at a time. For a unigrid Enzo simulation (i.e. no refinement), the minimum inter-particle separation in which gravity acts is twice the length of two cells, given as:
$\Delta x_{\text{Unigrid}}=2\frac{L}{N_{\text{DM}}^{1/3}}.$ (1)
When running Enzo simulations, we used a maximum refinement level of 8, meaning that the resolution was increased by a factor 28 in regions of high particle density. This means the highest resolution is given as:
$\Delta x_{\text{Enzo}}=\frac{1}{2^{8}}\Delta x_{\text{Unigrid}}=\frac{1}{2^{7}}\frac{L}{N_{\text{DM}}^{1/3}}.$ (2)
For the dark matter-only simulations carried out here, refinement is triggered once the particle over-density reaches a factor of 4 greater than the mean density (i.e. MinimumOverDensityForRefinement = 4). In addition to this we set the MinimumMassForRefinementLevelExponent = -0.1 which makes the refinement scheme super-Lagrangian and allows for the higher levels of refinement to be more easily triggered. The combination of these parameter choices mean that our simulation setup is somewhat conservative, employing an aggressive refinement strategy (see e.g. O’Shea et al., 2005).
@@
$\Delta x_{\text{SWIFT}}=\frac{1}{5^{2}}\frac{L}{N_{\text{DM}}^{1/3}}.$ (3)
@@ 2.2 Numerical Halo Finders
We analyse the simulation snapshots from our full suite of outputs at $z=20.0$, $z=15.0$ and $z=10.0$ using a friends-of-friends (FOF) (Efstathiou et al. (1985)) and HOP halo (Eisenstein & Hut (1998)) finder (results from the HOP finder can be found in the appendix of the paper as the results are very similar between FOF and HOP and the goal of this study is not to compare halo finders). The FOF halo finder accounts for the distances between dark matter particles within a single snapshot. We use a linking length $\Delta x(b)=b\times L/N_{\text{DM}}^{1/3}$, where $L/N_{\text{DM}}^{1/3}$ is the mean inter-particle separation. This refers to the maximum permitted separation between two particles. A group i.e. halo is found from a set of inter-linked particles.
@@
The HOP halo finder accounts for the distances between particles as well as the computed density of each particle. Rather than creating a continuous density field in the simulation volume, the density of each particle at its given position is estimated using a spherically symmetric cubic spline kernel (Monaghan & Lattanzio (1985)) on its $N_{\text{dens}}=64$ nearest neighbours. The density is normalised to the average density of particles within the simulation box e.g. a particle with a computed density of $\delta=80$ refers to a position 80 times denser than the average i.e. an overdensity (Skory et al. (2010)). A link is made by hopping from a given particle to the densest of its $N_{\text{hop}}=64$ neighbours. This process continues, forming a chain of increasing particle density until we reach a particle that is its own densest neighbour. All chains sharing the same densest particle are part of the same group. If the maximum particle density within a group is above a set threshold overdensity $\delta_{\text{peak}}$, that group is defined as a halo. We use a threshold overdensity of $\delta_{\text{peak}}=100$. This value was chosen to calibrate the HOP finder with the FOF finder so both halo finders found a similar halo number and mass range for the Enzo , $L=0.5$ cMpc/h, $N_{\text{DM}}=512^{3}$, $z=10$ snapshot.
The halo masses found from these halo catalogues are the FOF and HOP masses i.e. the mass of the halo is defined as the sum of the masses of the particles that make up the halos. After we create halo catalogues based on these halo finders, we only include halos made up of at least $N_{\text{DM, min}}=100$ particles to reduce numerical over-counting (i.e. halos must be well resolved (by at least 100 particles) before we identify them as halos).
We find that there is little difference between the number densities found using FOF and HOP for a given simulation suite, with HOP underestimating the FOF number densities by a factor of $\approx$ 2 at most (see Figure 1 for a comparison). Given that the halo finder parameters were calibrated at $z=10.0$, it is not surprising that there is little mass variance in their number density ratio at this redshift. From this point onwards we show the results for FOF only and leave those from HOP to the appendix.
Figure 1: Comparing the FOF and HOP halo number densities derived from Enzo and SWIFT simulation data (with halo masses on a logarithmic scale). We show how the ratio of the HOP number density to the FOF number density varies with halo mass. The HOP number density is within a factor of 2 of its FOF counterpart - particularly at the z = 10 outputs. Some larger deviations at higher z are seen as expected.
@@ 2.3 Numerical Halo Number Densities
A halo catalogue is created from each snapshot with a specific halo finder using the yt [^1] We use a version of yt called yt-swift developed by Rennehan (2022) as the most recently available version of yt cannot load SWIFT output data. and yt-astro-analysis packages (Turk et al. (2011), Smith et al. (2022)). An array of halo mass values is generated from each halo catalogue, filtered to exclude halos consisting of $<\,N_{\text{DM, min}}$ particles. For each array, the halo mass values are binned on a logarithmic scale into 24 histogram bins with $M_{\text{min}}=10^{3.75}\,\text{M}_{\odot}/\text{h}$ and $M_{\text{max}}=10^{11.75}\,\text{M}_{\odot}/\text{h}$. Halo catalogues from a given simulation suite, halo finder and redshift, $z$, are combined, giving rise to 12 datasets (see Table 3 for more detail). For each Enzo dataset, we combine halo catalogues based on 12 simulation boxes. For each SWIFT dataset, we combine halo catalogues based on 13 simulation boxes (see Tables 1 and 2 for more detail). In our algorithm below, we index datasets with $i\in\{1,\cdots,12\}$, mass bins with $j\in\{1,\cdots,24\}$, Enzo simulation boxes with $k\in\{1,\cdots,12\}$, and SWIFT simulation boxes with $k\in\{1,\cdots,13\}$.
@@
Table 3: The 12 datasets derived by combining halo catalogues for each simulation, halo finder, and redshift combination.
@@
We compare our numerical halo number densities with those derived from both analytical forms and popular fits. The number density (units: $\text{h}^{3}\,\text{cMpc}^{-3}$) of dark matter halos at a given redshift, $z$, in a mass bin centred at $\log_{10}(M_{\text{mid}}/\text{h}^{-1}\,\text{M}_{\odot})$ with a width of $\Delta\log_{10}M$ is defined as:
@@
The halo mass function (units: $\text{h}^{4}\,\text{cMpc}^{-3}\,\text{M}_{\odot}^{-1}$) is the differential halo number density per unit mass and is defined as:
@@
Here $\rho_{0}$ (units: $\text{h}^{2}\,\text{M}_{\odot}\,\text{cMpc}^{-3}$) and $\sigma(z,M)$ refer to the mean density of the universe and mass variance respectively. The mass variance is given as:
@@
Figure 2: Comparing the Enzo and SWIFT halo number densities with five halo number densities derived from fits (see Table 4) at $z=20.0$ (upper panel), $z=15.0$ (centre panel) and $z=10.0$ (lower panel) using the FOF halo finder (with halo masses and number densities on a logarithmic scale). The black rectangles indicate the regions for more detailed analysis as seen in Figure 3. The yellow shaded regions represent the approximate range of halo masses detected by JWST at $z\geq 10.0$. Over the halo mass range selected at $z=10.0$, all but the PS fit agree with numerical results within a factor of 2.
Figure 3: A zoom-in onto the black rectangles identified in Figure 2 (with halo masses and number densities on a logarithmic scale). These mass ranges represent the most massive halos accessible via numerical simulation at these redshifts. At this higher level of detail we see some discrepancy between the numerical halo mass functions of Enzo and SWIFT and their (semi-)analytical counterparts, particularly at $z=15.0$. The residuals for each of the lines are shown in Figure 4 and Figure 5.